# Token-Aware Coding Flow: A Study with Nano Surge in Reasoning Model thanks: * Corresponding authors

### v1

@@ IV Experimental Results
IV-A RQ2: Impact of Code Refactoring on Token Consumption under Functional Consistency Constraint
@@
To answer RQ2, we designed an experiment to investigate how code refactoring—specifically the elimination of code smells—affects token consumption during model inference. We first used the DeepSeek‑R1 inference model on both smelly_code.jsonl and clean_code.jsonl to establish baseline token‑usage data. Next, we applied automated refactoring to the smelly code samples, removing common smells (e.g., long methods, duplicated logic) to generate a refactored code set (rf_code). We then re‑ran inference on these refactored samples and compared their token consumption to the original smelly code. To ensure functional consistency, we employed CodeBLEU and other metrics to verify that the refactored code preserved the original behavior.
@@
During analysis, we recorded both the inference duration and token consumption for each trial and computed token consumption per unit time. Under the functional consistency constraint, refactored code is expected to be leaner, reducing unnecessary complexity and redundant patterns. Our results show that after refactoring, the DeepSeek‑R1 model performs inference more efficiently: despite maintaining equivalent functionality(Tab. III), the streamlined code requires significantly fewer tokens(Fig. 6)(Tab. II). In particular, refactoring complex smells—such as excessively long functions and duplicated segments—substantially reduced the number of reasoning steps and validations the model must perform.
@@
IV-B RQ3: Impact of Different Types of Code Smells on Token Consumption
@@
The first graph(Fig. 7) and the table(Tab. IV))present the average growth rate of token consumption across various code smell types. It is evident that most of the code smells, such as binary_operator_in_name, complicated_boolean_expression, and complicated_regex_expression, lead to a significant increase in token consumption. These types of smells are often characterized by complex logical constructs that demand deeper reasoning from the model, resulting in more tokens being processed[26]. Conversely, the smells func_name and too_long_parameter_list have a relatively lower impact on token growth. This suggests that these types of smells are simpler and less likely to introduce complexity in reasoning, causing fewer tokens to be consumed.
The second graph(Fig. 8) categorizes code smells into four groups: Naming, Expression, Structure, and Design. The Expression category stands out with the highest average token consumption growth, followed by Structure. These categories typically involve complex logical constructs or convoluted code structures, which require more tokens for the model to parse and reason about. In contrast, the Naming and Design categories exhibit lower token consumption growth. Issues like poor naming conventions and design flaws might affect code readability and maintainability, but they do not significantly complicate the reasoning process, leading to lower token usage by the model.
@@
IV-C RQ4: Impact of Explicitly Indicating Code Smell Types in Prompts on Token Consumption[27]
@@
IV-D RQ5: Prompt Engineering Strategies beyond Refactoring
@@
All experiments were performed on the same dataset of 300 smelly code snippets (smelly_code.jsonl). For each condition, we invoked DeepSeek‑R1, recording the total token consumption and calculating token consumption per unit time. Results were compared against a baseline prompt, which did not include any specific optimization strategies.
@@
Figure 10: Nano Surge on Token Consumption
Findings – Prompt Optimization Strategies Context Awareness and Responsibility Tuning emerged as the most effective strategies for reducing token consumption beyond refactoring(Tab. IX). Context Awareness decreased token usage by 15–20%, especially on lengthy or complex code. Responsibility Tuning achieved a 10–15% reduction by framing the model’s reasoning role. Cost Sensitivity, while reducing tokens by 20–30%, risked incomplete code generation under strict limits. These results suggest that combining context-rich prompts with precise role framing offers substantial efficiency gains without compromising functionality. Furthermore, the optimizations applied through these strategies also led to improvements in code quality metrics. Specifically, the codeBLEU score increased by 38%, and function similarity was enhanced by 4%, compared to the baseline performance. These improvements indicate that not only were token consumption and reasoning efficiency optimized, but the overall quality of the generated code also saw significant gains, further validating the efficacy of these prompt engineering strategies.
V Related Work
@@
V-A Code Smell Detection and Refactoring
@@
V-B LLM‑based Code Generation and Optimization
@@
V-C Integration of Code Quality and Inference Efficiency
@@
VI Conclusion and Future Work

### v2

@@ IV Experimental Study: Triggering the nano surge
IV-B RQ2: Impact of Code Refactoring on Token Consumption under Functional Consistency Constraint
@@
To address RQ2, we examined how refactoring to remove code smells impacts token consumption during model inference. Using DeepSeek‑R1, we first established baseline token usage on both smelly_code.jsonl and clean_code.jsonl. We then applied automated refactoring to the smelly code (removing common smells like long methods and duplicated logic) to produce a refactored code set (rf_code), and measured token usage on these samples. Functional consistency was verified using CodeBLEU and related metrics.
@@
For each run, we tracked inference duration and token consumption, and normalized by time. Results show that refactored code leads to more efficient inference—maintaining functionality (Tab. III), but requiring substantially fewer tokens (Fig. 7, Tab. II). Refactoring complex smells, especially long functions and duplicate code, significantly reduced the number of reasoning steps and validations the model performed.
@@
IV-C RQ3: Impact of Different Types of Code Smells on Token Consumption
@@
The first graph (Fig. 8) and table (Tab. IV) present the average growth rate of token consumption for different code smell types. Most smells, such as binary operator in name, complicated boolean expression, and complicated regex expression, cause a notable increase in token usage, likely due to their complex logic requiring deeper model reasoning [22]. In contrast, smells like func name and too long parameter list have a lower impact, indicating they are less complex and lead to only minor increases in token consumption.
The second graph (Fig. 9) groups code smells into Naming, Expression, Structure, and Design. Expression-related smells show the highest token consumption growth, followed by Structure. These categories often involve intricate logic or convoluted code structures, demanding more tokens for the model to process. By comparison, Naming and Design smells have a smaller effect on token growth, as they are less likely to complicate the reasoning process.
@@
IV-D RQ4: Impact of Explicitly Indicating Code Smell Types in Prompts on Token Consumption[23]
@@
IV-E RQ5: Prompt Engineering Strategies beyond Refactoring
@@
All experiments were performed on the same dataset of 300 smelly code snippets. For each condition, we invoked DeepSeek‑R1, recording the total token consumption and calculating token consumption per unit time. Results were compared against a baseline prompt, which did not include any specific optimization strategies.
@@
Findings – Prompt Optimization Strategies Context Awareness and Responsibility Tuning proved most effective for reducing token consumption beyond refactoring (Tab. IX). Context Awareness reduced token usage by 15–20%, especially for complex code, while Responsibility Tuning achieved a 10–15% reduction by specifying the model’s reasoning role. Cost Sensitivity reduced tokens by up to 30% but sometimes caused incomplete outputs. Overall, combining context-rich prompts with clear role framing offered notable efficiency gains without sacrificing functionality. These optimizations also improved code quality: codeBLEU scores increased by 38%, and function similarity improved by 4% compared to the baseline, demonstrating that prompt engineering can simultaneously enhance efficiency and output quality.
V Discussion
This study has conducted preliminary observations and analyses of the token consumption behavior exhibited by Large Language Models (LLMs) in Chain of Thought (CoT) scenarios[4]. The findings reveal complex interactions between code characteristics and token overhead, alongside several possibilities for optimization through external strategies[24]. The specific behavior of LLMs in software engineering tasks, particularly their resource consumption patterns (such as token usage), often demonstrates a high degree of contextual dependency and certain ”black box” characteristics, rendering precise prediction solely through theoretical models challenging.
In this context, the long-standing empirical tradition within software engineering—which emphasizes understanding and mastering complex systems through iteration, observation, experimentation, and data-driven analysis—offers a potent perspective and methodology for addressing such challenges[25]. When confronting the specific issue of token consumption by LLMs in code processing, an approach grounded in empiricism is particularly crucial[26]. Empirically observing the actual impact of various factors, such as different code features and prompting strategies, on token consumption, and quantitatively evaluating the effectiveness of intervention measures (like the code refactoring and prompt optimization explored in this study), serves as an effective pathway to deeply understand and manage this novel form of ”resource consumption.” The analytical framework of the present study represents a preliminary endeavor guided by this philosophy.
Looking ahead, the role of LLMs in software development and associated areas of focus are expected to continually evolve. On one hand, as the capabilities of LLMs iteratively improve, the conformity of the code they generate, or their ability to understand and intelligently refactor existing code[12], is poised for significant enhancement. This may gradually reduce the future need for manual attention to, and remediation of, traditional ”code smells,” potentially shifting the focus of software quality assurance to new dimensions. On the other hand, even if LLMs can efficiently process or generate ”smell-free,” high-quality code, the reasoning and generation processes essential for fulfilling complex requirements will inherently continue to incur token costs[27]. Therefore, regardless of how the issue of ”smells” at the code level evolves, attention to token consumption will remain a core concern in the field of LLM-assisted software development. Optimizing the token efficiency of the generation process and striking the optimal balance between output quality, functional completeness, and resource expenditure will continue to be key challenges and top priorities for future research aimed at the sustainable application of this technology[28].
VI Related Work
@@
VI-A Code Smell Detection and Refactoring
@@
VI-B LLM‑based Code Generation and Optimization
@@
VI-C Integration of Code Quality and Inference Efficiency
@@
VII Conclusion and Future Work